OpenStudy (anonymous):

Marsha is buying plants and soil for her garden. The soil cost $4 per bag, and the plants cost $10 each. She wants to buy at least 5 plants and can spend no more than $100. Write a system of linear inequalities to model the situation.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

show me what you have so far

OpenStudy (anonymous):

4x+10y<=100

OpenStudy (anonymous):

I am not sure how to write the second equation.

jimthompson5910 (jim_thompson5910):

"The soil cost $4 per bag, and the plants cost $10 each" "can spend no more than $100" that info definitely leads to `4x+10y<=100`. You are correct

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

x = number of bags of soil bought y = number of plants bought

jimthompson5910 (jim_thompson5910):

" She wants to buy at least 5 plants" so `y >= 5` we want the value of y to be 5 or greater

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

so the system is \[\Large \begin{cases}4x+10y\le100\\y\ge5\\x\ge0\end{cases}\] we have `x>=0` in there to ensure x is not negative

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thanks for your assistance! I need help with 2 more if that's fine with you.

jimthompson5910 (jim_thompson5910):

sure just post as a new question